OpenStudy (ayeeeitskateee):

For each function, what is the output of the given input? For f(x) = 4x + 7, find f(4) Please show work and give me an explanation. I am trying to study for my test with these practice problems. I'm not sure how to do this one. Please explain guys!! Will medal/fan as always.

OpenStudy (freckles):

f(x) is similar to f(4) but instead of x you see 4 so instead of the input being x the input is now 4 just replace the x's with 4's in the following: f(x)=4x+7 so you have f(4)=4(4)+7

OpenStudy (calculusxy):

our input is with f(x) and whatever answer to the function that was given with f(x) would be the output... in other words, if we have \[f(x) = 3x + 1 \] and we are said to find f(2), what we have to do is replace the x in the function with 2. \[f(2) = 3(2) + 1 \] simply solve \[3(2) + 1 = 6 + 1 = 7 \]

OpenStudy (calculusxy):

use that same principle with your problem, replace x with 4 \[\large f(4) = 4(4) + 7\] find the sum of \[4(4) + 7 \] and that will be your output

OpenStudy (calculusxy):

did you get your answer ?

OpenStudy (ayeeeitskateee):

Hey guys. Just now got you answer. Let me read! :) hold on.

OpenStudy (ayeeeitskateee):

Thank you both! I wish I could give you both medals. :( Maybe you could medal each other? lol.